# Averaging of quantum channels via channel-state duality

### v1

@@ 5 Channel $t$-designs
Several notions of designs for channels have been proposed in the literature, often by pushing forward a design on a larger unitary group through a fixed Stinespring dilation or by imposing moment-matching conditions on ensembles of CPTP maps, see e.g. [26]. The channel $t$-design notion in Eq. (43) is tailored to the present twirling problem: it is the minimal finite condition required to reproduce the averaged Choi operator for collective pre/post actions, and it follows directly from the partial-transpose reduction and any weighted group $t$-design for $t=t_{\rm in}+t_{\rm out}$. This perspective complements other channel-design frameworks by making the symmetry and the induced representation on the Choi space explicit.
@@ 6 Conclusions and outlook
We developed a representation-theoretic framework for averaging (twirling) of quantum channels that is driven by channel–state duality. Starting from the definition of channel twirling as uniform pre- and post-processing by the symmetry group representations on the input and output spaces, we derived an explicit Choi-level formulation in which the twirl becomes a group average acting directly on the Choi operator under the induced representation $\pi^{\mathrm{out}}\otimes\bar{\pi}^{\mathrm{in}}$. This viewpoint identifies the twirl as a projection onto the commutant algebra and leads to closed-form expressions in terms of generalized Schur operator bases obtained from generalized Schur-Weyl duality.
@@
A central simplification arises in the collective setting $\pi^{\mathrm{in}}(U)=U^{\otimes t_{\mathrm{in}}}$ and $\pi^{\mathrm{out}}(U)=U^{\otimes t_{\mathrm{out}}}$, where the commutant of $U^{\otimes t_{\mathrm{out}}}\otimes\bar{U}^{\otimes t_{\mathrm{in}}}$ is the walled Brauer algebra. By exploiting a partial-transpose identity, we showed that channel twirling can be reduced to an ordinary Schur-Weyl twirl of the partially transposed Choi operator under $U^{\otimes(t_{\mathrm{in}}+t_{\mathrm{out}})}$. This removes the contragredient action at the cost of a final partial transpose and enables practical evaluation using permutation-operator technology rather than mixed Schur transforms or explicit walled Brauer idempotents.
We further extended the construction beyond compact unitary symmetries to averaging over non-unitary operations described by reductive groups via Cartan decomposition. The resulting twirl admits a decomposition into invariant-sector projections weighted by coefficients $\beta_{k}$ determined solely by the Abelian Cartan component, and an analogous partial-transpose reduction applies to avoid contragredient representations on the compact part. In addition, we discussed a dual averaging realization in which the (Cartan) twirl can be implemented as a probabilistic mixture of unitary-$1$-design channels acting on invariant sectors through embedded unitary operator bases. Finally, motivated by unitary designs, we introduced a notion of channel $t$-design for collective actions: any (weighted) group $t$-design for $t=t_{\mathrm{in}}+t_{\mathrm{out}}$ induces a finite reconstruction of the averaged Choi operator, in this case preserving the structure of the original symmetry group action.
Several directions follow naturally. On the constructive side, it is important to develop efficient exact and approximate finite averaging sets for broader classes of symmetry groups and representations, and to quantify the error incurred when replacing Haar/Cartan integrals by designs in the induced channel-$t$ setting. On the algorithmic side, the partial-transpose reduction suggests scalable implementations of channel twirls based on permutation-operator decompositions and Schur-Weyl projectors, with complexity controlled by $t_{\mathrm{in}}+t_{\mathrm{out}}$ rather than the full walled Brauer machinery. Finally, extending the present approach to approximate symmetries, to non-identical local dimensions, and to other physically motivated group actions (beyond collective tensor powers) may further broaden the range of channel-averaging problems that admit analytic reduction and finite, design-like realizations.

### v2

@@ 5 Channel $t$-designs
Several notions of designs for channels have been proposed in the literature, often by pushing forward a design on a larger unitary group through a fixed Stinespring dilation or by imposing moment-matching conditions on ensembles of CPTP maps, see e.g. [43]. The channel $t$-design notion in Eq. (43) is tailored to the present twirling problem: it is the minimal finite condition required to reproduce the averaged Choi operator for collective pre/post actions, and it follows directly from the partial-transpose trick and any weighted group $t$-design for $t=t_{\rm in}+t_{\rm out}$. This perspective complements other channel-design frameworks by making the symmetry and the induced representation on the Choi space explicit.
@@ 6 Conclusions and outlook
We have shown that channel twirling is most naturally and generally formulated at the level of Choi operators, where it becomes a genuine group twirl under the induced output–input representation. From this perspective, channel averaging is no longer merely an operational pre- and post-processing prescription, but a representation-theoretic projection acting directly on the Choi operator space. This shift of viewpoint provides a unified and explicit description of channel twirling that applies to arbitrary input and output representations and does not rely on case-by-case constructions.
While symmetry-restricted quantum channels—covariant or equivariant—have been extensively studied, existing works typically focus on characterizing the set of invariant channels rather than on the twirling map itself. In contrast, the present work treats the twirl as an explicit linear superoperator acting on Choi operators. Concretely, the Haar (or Cartan) average defines a completely positive linear map whose range coincides with the commutant of the induced representation, so that the twirl realizes a projection onto this invariant subspace in the operator-algebraic sense. This distinction is crucial: instead of merely identifying the structural constraints that invariant channels must satisfy, we obtain a constructive formula for the averaging map itself, allowing one to compute the averaged channel directly rather than inferring its form indirectly from symmetry considerations.
@@
In the collective setting, where the symmetry acts as tensor powers on multiple input and output subsystems, the relevant commutant is the walled Brauer algebra. A central insight of this work is that the technical complexity associated with mixed Schur–Weyl duality can be bypassed by a partial-transpose reduction: the channel twirl is mapped to an ordinary Schur–Weyl twirl of the partially transposed Choi operator under a single tensor-power representation. This reduction replaces contragredient actions by a final partial transpose and enables practical evaluation using standard permutation-operator techniques, without constructing explicit walled Brauer idempotents or mixed Schur transforms.
We further extended the Choi-level formulation of channel twirling beyond compact symmetries acting via unitary representations to reductive, generally non-compact groups via Cartan decomposition. In this setting, where uniform averaging is not available, the twirl decomposes into invariant-sector projections determined by the maximal compact subgroup, weighted by coefficients that depend solely on the Abelian Cartan component. To our knowledge, this provides the first systematic and constructive treatment of channel twirling for non-compact transformations at the Choi level, clarifying both its algebraic structure and its normalization properties.
Two finite realizations of channel averaging were also developed. First, we showed that the Cartan twirl admits a dual representation as a probabilistic mixture of unitary-$1$-design channels acting on invariant sectors through embedded operator bases. Second, motivated by unitary designs, we introduced a notion of channel $t$-design tailored to the present setting: any weighted group $t$-design with $t=t_{\mathrm{in}}+t_{\mathrm{out}}$ induces an exact finite reconstruction of the averaged Choi operator. Together, these constructions demonstrate that the abstract Choi-level twirl admits finite, symmetry-respecting implementations without leaving the representation-theoretic framework.
Several directions follow naturally from this work. On the constructive side, it is important to develop efficient exact and approximate finite averaging sets for broader classes of symmetry groups and representations. A key open problem is to quantify the approximation error incurred when replacing Haar or Cartan twirling by a finite weighted sum (design-like implementation). Concretely, one would like bounds on the deviation of the resulting twirling superoperator from the ideal one, either at the Choi level (e.g. in trace or Hilbert–Schmidt norm for Choi operators) or directly at the channel level (e.g. in diamond norm), with explicit dependence on $t_{\mathrm{in}}+t_{\mathrm{out}}$, the local dimension, and the quality/size of the design.
From an algorithmic perspective, the partial-transpose reduction suggests scalable implementations of channel twirls based on permutation operators and Schur–Weyl projectors, with complexity controlled by $t_{\mathrm{in}}+t_{\mathrm{out}}$ rather than by the full mixed Schur–Weyl machinery. More broadly, treating channel symmetries directly at the Choi level appears essential for extending symmetry-based reductions to approximate symmetries, non-identical local dimensions, and learning problems involving structured quantum noise.